Please help me find the slope intercept of the one passing through the point (-4,0) and parallel to the line y=3x+2.
Answer by TutorDelphia(193)

y=mx+b
Parallel lines have the same slope so our slope will be 3
So far we know
y=(3)x+b
We can sub in -4 and 0 for x and y and solve for b
0=(3)*(-4)+b
0=-12+b
b=12
now going back to y=mx+b
we know that m=(3) and b=12
y=(3)x-+12 is our parallel line